An isolated system consists of two objects of the same mass. One object has an initial temperature of 0 °C and its specific heat is c1. The other has an initial temperature of 100°C and its specific heat is c2=5 c1. What is the equilibrium temperature of the system, assuming that no phase changes take place for either object?
Answer #1

As there is no phase change taking place, when both are in contact with each other one will lose heat and one will gain heat. let us assume that final temperature is Tf.

so m * c1 * ( Tf - 0 ) = m * 5 * c1 * ( 100 - Tf )

Tf = 5 * ( 100 - Tf )

Tf = 83.33 C

feel free to drop comment if you have any doubt.